OpenStudy (anonymous):

A storeowner purchases 30 boxes of flash drives. Of those boxes, \frac{1}{5} are given to the employees, and the rest are placed on the shelves to be sold. Each employee receives \frac{2}{3} of a box. How many employees are there? 4 employees 9 employees 12 employees 15 employees

OpenStudy (texaschic101):

30 total boxes......1/5 are given to employees 1/5 * 30 = 6.....so 6 are given to employees.....leaving (30 - 6) = 24 on the shelf. each employee receives 2/3 of a box..and remember, employees were given 6 boxes 2/3 * 6 = 12/3 = 4.....so there are 4 employees